OpenStudy (anonymous):

Determine the vector equation of each line: a) Direction vector [1,-2] and through P(-4,11) b) Normal vector of [5,-2] and through P (-1,2) c) Parallel to x + 6y = 1 and through P (2, -5) d) Parallel to the z-axis and through P(-4,3,7) e) Parallel to [x,y,z] = [1,3,-2] = t[5,2,1] and through P(2,0,1)

OpenStudy (anonymous):

\[r=i-2j+\lambda \left( -4i+11j \right)\]

OpenStudy (anonymous):

That is just for a) correct? But thank you!

OpenStudy (anonymous):

eq of any line parallel to x+6y=1 is x+6y=c it passes through (2,-5) 2+6*-5=c c=2-30=-28 hence x+6y=-28

OpenStudy (anonymous):

\[\left\{ r-\left( -i+2j \right) \right\}.\left( 5i-2j \right)=0\] \[r.\left( 5i-2j \right)+5+4=0\] \[r.\left( 5i-2j \right)+9=0\]

OpenStudy (anonymous):

For x+6y=c is the answer simply x+6y=-28 for a vector equation?

OpenStudy (anonymous):

rest in the morning .good night.

OpenStudy (anonymous):

thank you so much for your help!